OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 3 7x - 5y = -11

OpenStudy (anonymous):

FIrst find the least common multiple of 7 and 3. Which is 21. So, for 3x-6y=3 multiply the equation by 7. For 7x-5y=-11 multiply the equation by 3. Now decide which equation you want to contain -21.

OpenStudy (anonymous):

okay so it would be -3

OpenStudy (anonymous):

Okay so multiply that equation by 7.

OpenStudy (anonymous):

YOu should get -21x + 42y = -21

OpenStudy (anonymous):

Then multiply the second equation by 3.

OpenStudy (anonymous):

You should get 21x-15y=-33

OpenStudy (anonymous):

Now you have (1) -21x+42y=-21 (2) 21x-15y=-33 Then you can eliminate the 21's. Then combine the y's

OpenStudy (anonymous):

So you have 27y=-54 which is y=-2